OpenStudy (anonymous):

The length of a rectangular garden is 8ft more than twice its width. The area of the garden is 64 sq ft. Find the length and width of the garden. Please explain solving process.

OpenStudy (jlastino):

let x be the width the length is 8ft more than the width...how would you express this?

OpenStudy (anonymous):

(x+8) (x) = 64

OpenStudy (jlastino):

right :D now how do you solve for x?

OpenStudy (anonymous):

x^2+8x+64

OpenStudy (jlastino):

Hmm there's something wrong...you transposed 64 to the left side meaning its sign changes...can you do it again?

OpenStudy (anonymous):

-64

OpenStudy (jlastino):

yes now equate our equation to 0

OpenStudy (jlastino):

Do you need help?

OpenStudy (anonymous):

i am trying to find multiples of 64 that add 8x

OpenStudy (anonymous):

8x+x^2-64 = (x )(x )

OpenStudy (jlastino):

ok we made a huge mistake....the length is 8 more than twice the width :))

OpenStudy (jlastino):

(x) (2x+8) = 64 2x^2+8x=64 2x^2+8x-64 x^2+4x-32=0 there

OpenStudy (anonymous):

k

OpenStudy (jlastino):

i leave the rest to you then

OpenStudy (anonymous):

im stuck at (x+8) (x-4)

OpenStudy (jlastino):

equate those factors to 0. what of those two could be a possible measure of anything?

OpenStudy (anonymous):

-8 and +4?

OpenStudy (jlastino):

yeah so from those two which one can be a measurement?

OpenStudy (anonymous):

8*4=32

OpenStudy (jlastino):

umm that's not what i meant...is there such thing a -8cm?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ohh I forgot the 2x (-8 +4)

OpenStudy (jlastino):

so 4 would be the length of?

OpenStudy (jlastino):

2x (-8+4) that's still not it

OpenStudy (anonymous):

width

OpenStudy (jlastino):

yes and the length would be ?

OpenStudy (anonymous):

64/4=16

OpenStudy (jlastino):

correct :D

OpenStudy (anonymous):

Length = 16 Width = 4